OpenStudy (anonymous):

what is the first derivative of sin^2(3t^2) + 3e^(-t) - 5t^2 cos (2t^3)??

OpenStudy (anonymous):

yes, of course

OpenStudy (anonymous):

well the first one can be written like this\[(\sin(3t^2))^2\] and then you can use the power rule and chain rule. the second is just the chain rule and the third is the product rule with the chain rule.

OpenStudy (anonymous):

\[2(\sin(3t^2)(\cos(3t^2))(6t)\]

OpenStudy (anonymous):

so for \[3e ^{-t} = -3e ^{-t}\] right?

OpenStudy (jmark):

Given sin^2(3t^2) + 3e^(-t) - 5t^2 cos(2t^3) differentiate with respect to =>d/dt sin²(3t²)+d/dt 3e^(-t)-d/dt5t^2 cos (2t^3) => -10 t cos(2t^3)|+6t sin (6t²)+30t^4 sin (2t^3)-3e^-t Learn more about Concept of Standard Deviation at http://goo.gl/rGnw84

OpenStudy (anonymous):

aishh, many thankz @JMark , i understand now :D